### v4

# Watching the Watchers: Exposing Gender Disparities in Machine Translation Quality Estimation

Abstract
Quality estimation (QE)—the automatic assessment of translation quality—has recently become crucial across several stages of the translation pipeline, from data curation to training and decoding. While QE metrics have been optimized to align with human judgments, whether they encode social biases has been largely overlooked. Biased QE risks favoring certain demographic groups over others, e.g., by exacerbating gaps in visibility and usability. This paper defines and investigates gender bias of QE metrics and discusses its downstream implications for machine translation (MT). Experiments with state-of-the-art QE metrics across multiple domains, datasets, and languages reveal significant bias. When a human entity’s gender in the source is undisclosed, masculine-inflected translations score higher than feminine-inflected ones, and gender-neutral translations are penalized. Even when contextual cues disambiguate gender, using context-aware QE metrics leads to more errors in selecting the correct translation inflection for feminine referents than for masculine ones. Moreover, a biased QE metric affects data filtering and quality-aware decoding. Our findings underscore the need for a renewed focus on developing and evaluating QE metrics centered on gender.111We release code and artifacts at: https://github.com/deep-spin/gender-bias-qe-metrics.
[…]
Watching the Watchers:
Exposing Gender Disparities in Machine Translation Quality Estimation
Emmanouil Zaranis∗♣♠, Giuseppe Attanasio∗♣, Sweta Agrawal♣, André F.T. Martins♣♠♢ ♣Instituto de Telecomunicações, Lisbon, Portugal ♠Instituto Superior Técnico, Universidade de Lisboa, Portugal ♢Unbabel, Lisbon, Portugal emmanouil.zaranis@tecnico.ulisboa.pt
1 Introduction
Quality estimation—the automatic evaluation of machine-translated content without human-written references (Callison-Burch et al., 2012)—has gained increasing interest in the natural language processing (NLP) and machine translation communities. Recent work has focused on building QE metrics aligned with human quality judgments using neural encoders Rei et al. (2020); Perrella et al. (2022); Rei et al. (2023); Juraska et al. (2023); Guerreiro et al. (2024) or autoregressive language models Kocmi and Federmann (2023b); Fernandes et al. (2023); Kocmi and Federmann (2023a); Xu et al. (2023), and has explored their use throughout the translation pipeline, e.g., for data filtering (Peter et al., 2023; Alves et al., 2024), training (Ramos et al., 2024a; Yan et al., 2023; He et al., 2024), and improving output quality (Freitag et al., 2022; Fernandes et al., 2022; Farinhas et al., 2023; Ramos et al., 2024b; Vernikos and Popescu-Belis, 2024).
Evidence has shown QE scores effectively measure translation quality aspects like adequacy and fluency (Guerreiro et al., 2024). However, little is known about whether exogenous factors, e.g., conformity to social norms, play a role. If QE metrics favor normative language, they may penalize non-conforming expressions. Gender norms exemplify this issue—gender-based discrimination can amplify certain groups’ visibility in favor of others, leading to representational and allocative harms (Crawford, 2017). For instance, biased metrics may systematically retain masculine-inflected data in parallel corpus filtering while discarding equivalent-quality feminine translations. Similarly, when using QE to improve MT systems (Fernandes et al., 2022), such metrics could lead to systems that predominantly generate masculine translations, forcing women users to spend more time and effort correcting gender errors (Savoldi et al., 2024).
In this paper, we conceptualize and measure gender bias in QE metrics and its downstream effects. We draw two conditions to identify a biased metric. A QE metric exhibits gender bias if, when translating human entities, it systematically (i) assigns higher scores to one gender form over others when no explicit lexical cues disclose gender, or (ii) fails to detect the correct gender form even when lexical cues are provided, disproportionately across genders. We investigate bias across commonly used QE metrics through an extensive empirical study. We focus on out-of-English translation where the target language uses grammatical gender,222I.e., where gender for human referents is assigned on a semantic basis and reflected with morphological marks. a scenario frequently used in prior research on gender bias in MT (e.g., Vanmassenhove et al., 2018; Stanovsky et al., 2019; Saunders et al., 2020; Attanasio et al., 2023; Sant et al., 2024). We address both binary (i.e., contrasting masculine and feminine translations) and gender-neutral (i.e., contrasting gendered and gender-neutral forms) translation scenarios, and experiment with 11 state-of-the-art metrics (supervised and prompted) across two domains, three datasets, and eight target languages.
Our analysis reveals evidence of gender bias across metrics, languages, and experimental conditions. Masculine forms receive higher scores when the source gender is ambiguous. Moreover, metrics favor gendered over neutral expressions. Even with disambiguating context, metrics show higher error rates in identifying feminine gender forms. These biases significantly impact MT pipelines: feminine-inflected translations fail quality thresholds more frequently than masculine equivalents despite equal validity. During inference, QE-based hypothesis reranking amplifies gender bias in MT while improving translation quality. These findings call for extending QE metrics evaluation beyond human judgment alignment to consider gender equity.
2 Related Work
2.1 Evaluating MT Metrics
Significant efforts have been made to evaluate automatic metrics for MT. Most have focused on comparing sentence- or system-level correlations between a metric’s score and human judgements (Macháček and Bojar, 2014; Stanojević et al., 2015; Kocmi et al., 2021; Freitag et al., 2023; Deutsch et al., 2023; Thompson et al., 2024). Finer-grained analyses have discovered poor robustness to low-quality translations (Fomicheva and Specia, 2019), domain (Zouhar et al., 2024), or passive voice (Avramidis et al., 2023). Metrics have also been shown to produce imbalanced score distributions (Sun et al., 2020), and fail to distinguish discrepancies in numbers and named entities (Amrhein and Sennrich, 2022) or high-quality translations (Agrawal et al., 2024b). However, none of the studies have thus far evaluated gender bias exhibited by QE metrics.
2.2 Gender Bias in MT
Gender bias in MT systems—stemming from societal norms, model design, and deployment decisions—has been extensively documented (Savoldi et al., 2021). Prior work in NLP has primarily focused on measuring the extent of such bias in system outputs (e.g., Stanovsky et al., 2019; Cho et al., 2019; Hovy et al., 2020; Vanmassenhove and Monti, 2021; Levy et al., 2021; Lucy and Bamman, 2021; Sant et al., 2024; Piergentili et al., 2023; Lauscher et al., 2023; Lardelli et al., 2024; Stewart and Mihalcea, 2024), as well as on proposing mitigation strategies (e.g., Escudé Font and Costa-jussà, 2019; Costa-jussà and de Jorge, 2020; Saunders et al., 2020; Saunders and Byrne, 2020; Escolano et al., 2021; Lee et al., 2023; Garg et al., 2024). In contrast to this line of work, we shift the focus from the translation systems themselves to the task of quality estimation, a step that, typically, follows the MT. Further, we investigate whether bias in QE can propagate to or reinforce bias in MT outputs.
2.3 Social Biases in NLG Metrics
Several studies have investigated whether metrics used to evaluate natural language generation (NLG) exhibit social biases. Qiu et al. (2023) test whether n-gram- and model-based metrics, e.g. CLIPScore (Hessel et al., 2021) show gender bias in image captioning. Sun et al. (2022) quantify several social biases. They use English coreference resolution examples to measure the gender bias of several n-grams and model-based NLG metrics, including BERTScore (Zhang* et al., 2020), MoverScore (Zhao et al., 2019), and BLEURT (Sellam et al., 2020). Gao and Wan (2022) measure race and gender stereotypes using WEAT (Caliskan et al., 2017) and SEAT (May et al., 2019). However, no study targets QE metrics for MT or reference-free NLG metrics.
3 Measuring Gender Bias in QE
Several design choices arise when testing how neural QE metrics handle gender-related translations, including (i) source and target languages, (ii) domain, (iii) granularity (document, passage, or sentence-level MT), (iv) gender ambiguity in the source, and (v) gender-inflected words in the target. In this work, we experiment with out-of-English (En*) MT setups where the target language (*) is characterized by grammatical gender. In these languages, masculine and feminine genders are represented through word choices and inflection (e.g., different suffixes, as in Figure 1), while neo-pronouns, neo-morphemes, or rephrasing permit beyond-the-binary or neutral translation (Lauscher et al., 2023; Piergentili et al., 2023; Lardelli et al., 2024; Piergentili et al., 2024). Following prominent work on gender bias in MT (Currey et al., 2022; Rarrick et al., 2023), we focus on sentence-level translations from multiple domains and observe gender-inflected words referring to human entities. We challenge QE metrics with translation pairs where the entity’s gender in the source is ambiguous or contextual cues disambiguate it.
3.1 Formalizing Quality Estimation
For the remainder of the paper, we will discuss QE metrics that support quality assessment of translations Graham et al. (2016); Burchardt (2013). A QE metric is thus a function such that:
where is the English source, the translation to evaluate, and is the assessment score. Typically, is a neural function parameterized by , and a is bounded in a range with the edges indicating the poorest and highest quality.
QE Metrics
We test 11 QE metrics, varying model sizes and architectures. We include seven neural metrics that ranked highest in recent editions of the WMT QE Shared Task (Blain et al., 2023): CometKiwi 22 (Rei et al., 2022), CometKiwi 23 XL/XXL (Rei et al., 2023), xCOMET XL/XXL (Guerreiro et al., 2024), and MetricX 23 L/XL (Juraska et al., 2023). Following Kocmi and Federmann (2023b), we test the GEMBA-DA prompt with three decoder-only, instruction-tuned, open-weight language models—Mistral 7B (Jiang et al., 2023), Gemma 2 9B (Team et al., 2024), Llama 3.1 70B (Dubey et al., 2024)—and a commercial API model, GPT 4o (Achiam et al., 2023).
Our analysis will investigate whether metrics can use extra-sentential contextual cues (§3.2.3). However, contemporary QE metrics are not trained on nor support context beyond the sentence level off-the-shelf. Thus, we rely on two inference-time strategies to incorporate extra-sentential context Vernikos et al. (2022): 1) We concatenate the context string to and , i.e., , and 2) we apply the same strategy as (1) but automatically translate the context for the hypothesis, i.e., .333We use NLLB-200 3.3B (NLLB Team et al., 2022) for translating as it supports all studied languages. The latter variant prevents the creation of a cross-lingual hypothesis, out-of-distribution for the QE metric. We modify the prompt for GEMBA metrics to account for context (full details in Appendix A.1).
3.2 Experimental Conditions
In §1, we defined a biased QE metric as one that exhibits systematic undue preferences for gender forms in the target or disproportionate error rate across gender groups. Here, we describe the datasets and measurement methods we use to assess the presence and extent of gender bias.
3.2.1 Datasets with Minimal Edits
We study datasets organized in minimal edit contrastive pairs, where sources and reference translations differ only in gender-related words. This approach has two main advantages. First, it enables precise isolation of effects related to gender-inflected elements (see, for example, translations in Figure 1). Therefore, any change in the QE score can only be attributed to the metric’s sensitivity to those words. Second, using human-written references neutralizes potential confounding quality-related effects.444Sentence-level outputs are converging toward human-level translation quality in MT research; thus, using references is a valid choice. However, we relax this constraint in §5.2.
3.2.2 Gender-Ambiguous Entities
With gender ambiguity in the source, we set out to study two contrastive setups: masculine (m) vs. feminine (f) and gendered vs. gender-neutral translations. For the former (m/f) experiments, we use MT-GenEval (Currey et al., 2022) and GATE (Rarrick et al., 2023), two established sentence-level corpora for gender bias evaluation in MT with translations provided by professional translators. From MT-GenEval, we retrieve all test examples from the contextual subset, discarding the context that would disambiguate the gender. From GATE, we collect all instances with one gender-marked entity. The two corpora provide linguistic diversity: MT-GenEval collects sentences from Wikipedia, while sources in GATE are written by linguists and cover diverse linguistic phenomena. MT-GenEval includes eight target languages (ar, de, es, fr, hi, it, pt, ru) and GATE three (es, fr, it). Following prior work (Beutel et al., 2019; Gaut et al., 2020; Attanasio et al., 2024), we observe gaps in relativistic terms. We measure the ratio where and are the feminine- and masculine-inflected translations, respectively. Echoing the desiderata of counterfactual fairness (Czarnowska et al., 2021), the two scores should not differ, i.e., the ideal ratio is 1.
We use mGeNTE (Savoldi et al., 2025) for gender-neutral experiments. The benchmark targets gender-neutral translations of English sources written by professional translators for it, es, de.555Neutralization strategies might extend beyond simple edits. While breaking the minimal-edits requirement, these instances highlight meaningful contrasts between common gendered translations and less common neutral ones. We focus on ambiguous sources (Set-N), and measure , the ratio between the gender-neutral and gendered score. Since gender-neutral translations provide inclusive alternatives that maintain semantic accuracy, QE metrics should assign them higher scores than their gendered counterparts, i.e., the ratio should ideally exceed 1. However, equality between scores represents a minimum acceptable threshold.
3.2.3 Gender-Unambiguous Entities
We study cases where intra- or extra-sentential lexical cues disambiguate gender. For the intra-sentential case, we use the MT-GenEval’s counterfactual subset. It consists of paired instances with masculine- and feminine-inflected sources, each with its corresponding reference. For the extra-sentential case, we use the MT-GenEval’s contextual subset, where each test instance is paired with a preceding sentence disambiguating the gender. Figure 1 reports an example of the extra-sentential case. An instance with intra-sentential cues would be: “In her thirties, Tymoshenko has worked as a practicing economist and academic.”
Since a single correct gender form exists, given a set of instances , we compute the error rate ()—the fraction of instances where the QE metric assigns a higher score to the incorrect gender form. We compute it for all instances and separately for feminine and masculine sources, i.e., and , respectively.666We count ties as errors since they indicate the system’s failure to select the correct gender inflection. To measure gender disparities, we adopt a Multi-group Comparison Metric (Czarnowska et al., 2021) defined as:
Loosely inspired by False Positive Equality Difference (Dixon et al., 2018), the metric operationalizes the statement: “Even when lexical cues leave no doubts on gender identity, sources with feminine referents exhibit times the number of errors in gender inflection as sources with masculine referents.” The ideal scenario is equal errors, i.e., a ratio of 1. We report , i.e., the instance-level ratio of the QE scores assigned to feminine and masculine pairs, in Appendix B.2.
4 Findings
This section presents the results for ambiguous (§4.1), unambiguous with intra-sentential (§4.2.1) and extra-sentential cues (§4.2.2) pairs. Our analysis reveals significant () gender bias across most evaluated metrics and configurations (see Appendix A.2 for details on statistical testing).
| Metric | ||
| CometKiwi 22 | 0.11 | 1.70 |
| CometKiwi 23 XL | 0.09 | 1.18 |
| CometKiwi 23 XXL | 0.07 | 0.87 |
| xCOMET XL | 0.10 | 1.81 |
| xCOMET XXL | 0.08 | 1.32 |
| MetricX 23 L | 0.31 | 1.25 |
| MetricX 23 XL | 0.12 | 1.19 |
| Llama 3.1 70B | 0.31 | 1.16 |
| Gemma 2 9B | 0.28 | 1.36 |
| Mistral 7B | 0.74 | 1.13 |
| GPT 4o | 0.16 | 1.15 |
4.1 Gender-Ambiguous Instances
Most QE metrics and language pairs show a significantly higher score for masculine-inflected references. Figure 2 reports the average f/m ratio of QE scores on MT-GenEval test sets (full raw scores are in Figure 5 of Appendix B.1). We found no significant difference between f/m scores only for MetricX (X and XL), ende. The metric is generally the closest to parity across languages. We hypothesize that MetricX models, unlike Comet, are trained on additional synthetic data that attributes perfect scores to reference translations. This choice likely skews the metric results by making it more sensitive to identifying references than quality differences from minimal edits. Among the best models, Mistral 7B reaches an average ratio of 0.9961. Notably, preference toward the masculine-inflected translations increases with size for both CometKiwi (22 < 23 XL < 23 XXL) and GEMBA models (Mistral 7B < Gemma 2 9B < Llama 3.1 70B). These results hold for GATE (see Figures 6 and 7 in Appendix B.1).
Most QE metrics score gender-neutral translations consistently lower than gendered ones on all languages. Figure 8 in Appendix B.1 reports the ratio between QE scores for gendered and gender-neutral translations on mGeNTE’s ambiguous instances. Despite the unknown gender in English, all metrics but MetricX and Mistral 7B prefer gendered translations significantly. Similar to the f/m setup, we observe that the larger a CometKiwi variant is, so is the preference for gendered translations. Findings are consistent across languages, with a larger magnitude on enes, suggesting that QE metrics may still be unsuitable for assessing gender-fair forms Sczesny et al. (2016).
4.2 Gender-Unambiguous Instances
Tables 1 and 2 present the total error rate () and error rate ratio () for the intra and extra-sentential cases across metrics.
4.2.1 Intra-sentential Cues
Most metrics consistently exhibit a higher frequency of errors when the source mentions a feminine entity than a masculine one. As shown in Table 1, the majority of metrics yield values greater than 1 for , indicating increased susceptibility to errors when tasked to select the correctly inflected translation and the referent in the source is feminine. CometKiwi 23 XXL is the only exception, yielding the least biased behavior overall; however, it still exhibits bias in 2 out of 8 languages (full results disaggregated by language are in Table 8 of Appendix B.2).
4.2.2 Extra-sentential Cues
| Metric | ||||
| CometKiwi 22 | 0.28 | 0.84 | 0.19 | 6.68 |
| CometKiwi 23 XL | 0.28 | 0.92 | 0.17 | 5.09 |
| CometKiwi 23 XXL | 0.14 | 2.20 | 0.14 | 4.85 |
| xCOMET XL | 0.29 | 4.53 | 0.24 | 7.03 |
| xCOMET XXL | 0.20 | 4.97 | 0.20 | 5.47 |
| MetricX 23 L | 0.39 | 0.73 | 0.34 | 1.53 |
| MetricX 23 XL | 0.22 | 1.54 | 0.28 | 1.60 |
| Llama 3.1 70B | 0.61 | 2.20 | - | - |
| Gemma 2 9B | 0.57 | 2.71 | - | - |
| Mistral 7B | 0.88 | 1.22 | - | - |
| GPT 4o | 0.44 | 2.64 | - | - |
Most QE metrics exhibit a significantly higher error rate on feminine sources than masculine sources. When the context is not translated, four out of seven context-aware neural metric variants make at least 1.5 times as many errors on feminine sources as, on average, masculine ones. CometKiwi 23 XL yields the best error rate ratio among neural metrics, followed by CometKiwi 22 and MetricX 23 L.
Metrics that use translated context yield lower overall error rates. However, improved accuracy is primarily driven by better results on masculine sources. Therefore, translated context exacerbates gender disparities for all QE metrics (see full results in Table 10 of Appendix B.2). increases on average 3 times, with the highest increase accounting for CometKiwi 22 (approximately 8 times higher error rate).777We observed similar trends even when using Google Translate for translating (see Appendix B.2). We hypothesize that this significant increase might be attributed to the translated context, which may induce bias towards masculine referents. These results suggest that only looking into metrics optimizing the overall error rate can be misleading and result in higher disparities.
Regarding GEMBA metrics, only Mistral 7B shows a balanced error rate between genders. On the contrary, Llama 3.1, Gemma 2, and GPT-4o make approximately two to three times as many errors on feminine sources as on masculine sources. These findings indicate a strong preference for masculine-inflected translations, even when gender cues are available in the preceding context.
However, upon a closer inspection, we noted that GEMBA QE metrics cannot distinguish gender inflections successfully. Ties—i.e., an equal DA score for the correct and incorrect translation—are widespread. On average, Mistral 7B assigns the same score to 85%, Gemma 2 9B to 50%, Llama 3.1 70B to 55%, and GPT-4o to 38% of test samples (full details in Table 14, Appendix B.2). Even with fewer ties, (Gemma 2 and GPT-4o) metrics show a significantly () higher error rate for feminine sources. See Table 11 of Appendix B.2 for results on individual languages. We explore further GEMBA models’ brittleness to surface-level differences in the following section.
4.3 Bias and Surface-Level Brittleness
An unbiased QE metric evaluates masculine and feminine translations equivalently when the source gender is unknown. However, whether such scores derive from an unbiased assessment or failure to detect gender markers remains unresolved, which could compromise accuracy and usability when gender is known. To investigate this distinction, we analyze the metrics results jointly across ambiguous and unambiguous scenarios.
Figure 3 reports the total error rate when source gender is known—a measure of overall accuracy—vs. gender bias measurements in ambiguous () and unambiguous cases (). For simplicity, we report one minus these values to have the optimal result at 0. The chart reveals that among the metrics identified as least biased metrics in the ambiguous case (§4.1), only MetricX 23 XL can distinguish gender inflections in the unambiguous case (). In contrast, Mistral 7B scores many ties (Table 14 in Appendix B.2), demonstrating a low bias in both setups but an astonishingly high error rate (and extreme brittleness to gender-related surface-level differences). More generally, this shortcoming can be attributed to the tendency of GEMBA metrics to assign coarse-grained scores Zheng et al. (2024); Stureborg et al. (2024)—typically 85, 90, 95, and 100 in our experiments. Based on this result, we discourage using GEMBA-prompted language models if capturing gender-related phenomena is crucial for the use case at hand.
CometKiwi 23 XXL is a better alternative, being the only other Pareto-optimal metric in ambiguous and unambiguous setups with a low overall error rate. Comparing the xCOMET and Kiwi families, we note that they achieve similar gender error rates, but the latter is fairer, on average. This holistic view is critical, as it can inform the selection of metrics for downstream applications where accurate and unbiased evaluations are necessary.
5 Downstream Implications
Previous sections have collected evidence of widespread gender bias in contemporary QE metrics. This section explores the downstream implications of using biased QE metrics in different stages of a classical MT pipeline.
5.1 Quality Filtering
Filtering high-quality parallel data from web crawls is a long-standing challenge in MT research. Evolving from traditional approaches using heuristics (Resnik and Smith, 2003) or noise detection techniques (Taghipour et al., 2011), recent work has adopted QE metrics to filter large corpora (Peter et al., 2023; Alves et al., 2024). Arguably, data retained post filtering influences each subsequent step in the MT pipeline.
In §4.1, we show that in the absence of clear gender cues in the source, contemporary QE metrics tend to favor masculine-inflected hypotheses, casting shadows over their use as quality filters. To exemplify the issue, Figure 4 shows a complementary cumulative distribution function of a distribution of CometKiwi 23 XXL scores (Koenecke et al., 2020; Attanasio et al., 2024). The chart illustrates explicitly the number of instances that would be filtered out at a given quality threshold, separately by gender. A threshold of 0.8 would retain 75% of masculine instances but only 63% of feminine ones. This disparity highlights that even when quality is comparable, sentences with feminine forms are more likely to be filtered out, potentially amplifying gender representation imbalances.
5.2 Machine Translation Quality Assessment
| Metric | ||
| CometKiwi 22 | 0.13 | 1.34 |
| CometKiwi 23 XL | 0.11 | 1.60 |
| CometKiwi 23 XXL | 0.09 | 1.62 |
| xCOMET XL | 0.12 | 1.96 |
| xCOMET XXL | 0.10 | 1.23 |
| MetricX 23 L | 0.34 | 1.17 |
| MetricX 23 XL | 0.14 | 1.33 |
Our evaluation in §4 focuses on a controlled setup with minimally contrastive reference translations to isolate the impact of gender-inflected linguistic elements on QE metrics. However, QE metrics are primarily used to assess machine-generated translations—which often vary in style and quality (e.g., exhibiting characteristics of translationese). We extend our analysis to investigate whether these metrics demonstrate systematic biases when applied to automatic translations. We focus on cases where intra-sentential cues resolve the gender identity of the referent, as in §4.2.1. Specifically, we translate the source-feminine () and source-masculine () texts from the counterfactual subset of MTGenEval using Google Translate (GT).888Prior work has shown that GT is less biased when the context provides sufficient information to disambiguate the gender Rescigno et al. (2023); Piazzolla et al. (2023). To account for quality variations unrelated to gendered forms, we only retain the counterfactual pairs () and () where the translations and exhibit the correct gender inflection, as categorized by Currey et al. (2022), and are of similar quality as measured by BLEU Papineni et al. (2002).999Details about the filtering process are provided in Appendix A.3. Notably, approximately half of the initial instances are retained, as reported in Table 15 of Appendix A.3. Our findings reveal that QE metrics exhibit a significant bias when assessing automatic translations. When challenged to select the correctly inflected automatic translation, they are more prone to errors on sources with feminine referents than on those with masculine referents (Table 3).101010Individual language results are in Table 13 (App B.2). These results are in line with those we obtained from human-written references.
| Comet-22 | Match | |||
| Decoding | Metric | |||
| Greedy | - | 86.73 | 88.45 | -45.67 |
| QAD | CK 22 | 87.66 | 89.39 | -46.84 |
| CK 23 XXL | 87.46 | 89.11 | -43.47 | |
5.3 Quality-Aware Decoding
So far, we conceptualized and measured gender bias as an undue preference expressed by quality metrics. However, when used to guide the translation for improved quality, a biased metric can impact and potentially amplify gender bias of MT outputs. To investigate this intuition, we experiment with quality-aware decoding (Fernandes et al., 2022, QAD), a common downstream application of QE metrics. Unlike greedy decoding, which selects the most likely next token at each step, QAD approaches rank N-best translation hypotheses generated by an MT model using automatic metrics.
Setup
As the target translation system, we experiment with TowerInstruct-v0.2 Alves et al. (2024), a state-of-the-art model for MT. Using greedy and QAD, we generate translations of the MT-GenEval contextual (ambiguous) sources for six supported languages: de, es, fr, it, ru, and pt. For QAD, we generate 50 hypotheses using epsilon sampling () and perform reranking using CometKiwi 22 and CometKiwi 23 XXL. Based on the findings in §4.3, CometKiwi 23 XXL achieves Pareto-optimal performance and CometKiwi 22 exhibits bias in the ambiguous case, allowing us to compare the impact of differently biased metrics.
We evaluate translation quality using reference-based Comet-22 against masculine () and feminine () references. Relying on notions linking gender bias and uneven visibility (Savoldi et al., 2021), we measure bias in the MT system as the frequency each gender is represented in the generated hypothesis. Following Currey et al. (2022), we check if specific gender-marked words, uniquely present in either or , are present in the generated hypotheses. We report , the difference between the number of feminine and masculine reference matches. A close to zero indicates minimal gender bias as translations are equally matched to female and male references, while a large positive (negative) suggests that the system is biased towards female (masculine) inflections.
Findings
Table 4 presents the translation quality of outputs using both greedy and QAD methods. Using greedy decoding, TowerInstruct-v0.2 produces more masculine-inflected translations of higher quality, as evidenced by absolute higher Comet-22 scores for masculine outputs () and a large negative (-45.7). While reranking improves overall translation quality over greedy, as measured by Comet-22, for both gender groups, its effect on gender disparity varies depending on the metric used. Specifically, QAD with the CometKiwi 23 XXL model enhances fairness, achieving the best absolute values for ambiguous source texts. In contrast, using a biased metric such as CometKiwi 22, further exacerbates gender disparity, as indicated by the more negative values. Therefore, careful selection and calibration of QE metrics are crucial to ensure that improvements in overall translation quality do not inadvertently amplify gender bias.
6 Conclusion
We formalized and measured gender bias in quality estimation metrics commonly used to assess translation quality. We studied translation from English into languages with grammatical gender, isolating the gender lexical phenomenon via minimal contrastive pairs. Our findings revealed concerning patterns. Through extensive experiments across languages, domains, and datasets, we found most QE metrics systematically penalize feminine forms in the target language—both when gender is ambiguous and when it is contextually clear. Similarly, gendered forms are mostly preferred over gender-neutral forms. Moreover, we show that QE bias can affect data quality filtering and quality-aware decoding with MT systems.
We should, indeed, watch the watchers. Undue gendered preferences call for a new focus on how we develop and evaluate QE metrics. Future work should center gender representativeness in training data for QE metrics, including and normalizing gender-neutral instances. Moreover, new scrutiny of QE metrics entails moving beyond simple human alignment to create targeted challenge sets that can expose undue preferences.
Limitations
Our study is characterized by several methodological constraints that merit critical reflection. We limited our investigation to sentence-level MT, which, while representative of prevalent machine translation scenarios, does not cover phenomena potentially arising within more extended contexts, e.g., document-level Jiang et al. (2022); Raunak et al. (2023); Fernandes et al. (2025) and conversational MT Agrawal et al. (2024a). The English source sentences examined were derived from naturalistic passages of high quality and relatively limited syntactic variation.
Moreover, our evaluation of neural QE metrics with extra-sentential context, relies on inference-time strategies for integrating contextual information Vernikos et al. (2022); Agrawal et al. (2024a). We acknowledge that translating context may introduce additional bias depending on the specific translation system used. We leave a more in-depth investigation of this phenomenon—including more resource-intensive solutions, such as targeted fine-tuning or interpretability-driven strategies for selecting few-shot examples to mitigate biases (e.g., gender) in the translation of contextual information Thakur et al. (2023); Zaranis et al. (2024)—to future work.
Finally, methodologically, we implemented GEMBA-DA in a zero-shot configuration, aligning with contemporary research interested in pushing zero-shot problem-solving (Guo et al., 2025). We defer comprehensive exploration of few-shot and MQM-based prompting strategies (Kocmi and Federmann, 2023a) to future work.
Ethical Considerations
Using gender as a variable requires careful ethical considerations. The corpora we used, initially annotated for gender identity using textual markers like pronouns and titles, present two main concerns. First, gender is a nuanced, evolving attribute, making reliance on textual information alone potentially inaccurate, especially with unverifiable or outdated sources. Second, much of our analysis uses binary gender categories (masculine and feminine). However, we recognize gender as a spectrum and advocate for future studies to account for non-binary and other gender identities. To support gender inclusivity in machine translation, we assessed quality estimation on gender-neutral translations in Italian, Spanish, and German using the mGeNTE corpus.
Acknowledgements
We would like to thank Ben Peters, Dennis Fucci, António Farinhas, Sonal Sannigrahi for their feedback and insightful discussions in the earlier versions of the paper, and Anna Currey for providing valuable help with the MT-GenEval dataset. This work was supported by the Portuguese Recovery and Resilience Plan through project C645008882-00000055 (Center for Responsible AI), by EU’s Horizon Europe Research and Innovation Actions (UTTER, contract 101070631), by the project DECOLLAGE (ERC-2022-CoG 101088763), and by FCT/MECI through national funds and when applicable co-funded EU funds under UID/50008: Instituto de Telecomunicações.
Appendix A Experimental Details
A.1 Computing DA Scores
COMET.
We use the HuggingFace Hub for all COMET metrics and the official implementation from https://github.com/Unbabel/COMET. All metrics produce an inclusive score between 0 and 1, indicating the worst and best translation possible.
MetricX 23.
We use the official implementation found at https://github.com/google-research/metricx. Since these metrics produce error assessments, with a score between 0 (no errors) and 25 (maximum errors), we rescale the output as , where is the raw output of the metric.
GEMBA.
We use the same decoding configuration and post-processing for all the language models prompted with GEMBA-like prompts. Specifically, we use greedy decoding, with a maximum number of generated tokens of 256. Each model’s specific chat template is retrieved from the transformers’ implementation (Wolf et al., 2020). We use the instruction-tuned variants and a zero-shot prompt for all open-weight models. Non-contextual and context-aware prompts are reported in Table 6. We experimented two variants to include context in prompts, finding no significant difference among those. We finally used the v1 reported in Table 6.
[…]
The Hugging Face checkpoint and OpenAI’s model IDs are reported in Table 5.
| [1] | CometKiwi 22 | Unbabel/wmt22-cometkiwi-da |
| [2] | CometKiwi 23 XL | Unbabel/wmt23-cometkiwi-da-xl |
| [3] | CometKiwi 23 XXL | Unbabel/wmt23-cometkiwi-da-xxl |
| [4] | xCOMET XL | Unbabel/XCOMET-XL |
| [5] | xCOMET XXL | Unbabel/XCOMET-XXL |
| [6] | MetricX 23 L | google/metricx-23-qe-large-v2p0 |
| [7] | MetricX 23 XL | google/metricx-23-qe-xl-v2p0 |
| [8] | Mistral 7B | mistralai/Mistral-7B-Instruct-v0.2 |
| [9] | Gemma 2 9B | google/gemma-2-9b-it |
| [10] | Llama 3.1 70B | meta-llama/Meta-Llama-3.1-70B-Instruct |
| [11] | GPT 4o | gpt-4o-2024-05-13 |
| Format | Template |
| GEMBA |
Score the following translation from {src_lang} to {tgt_lang} on a continuous scale from 0 to 100, where score of zero means ‘‘no meaning preserved’’ and score of one hundred means ‘‘perfect meaning and grammar’’.
{src_lang} source: ‘‘{source}’’ {tgt_lang} translation: ‘‘{hyphothesis}’’ Score: |
| GEMBActx v1 |
Score the following translation from {src_lang} to {tgt_lang} on a continuous scale from 0 to 100 given the preceding context, where score of zero means ‘‘no meaning preserved’’ and score of one hundred means ‘‘perfect meaning and grammar’’.
{src_lang} source: ‘‘{context}’’ {src_lang} source: ‘‘{source}’’ {tgt_lang} translation: ‘‘{hyphothesis}’’ Score: |
| GEMBActx v2 |
Score the following translation from {src_lang} to {tgt_lang} on a continuous scale from 0 to 100, where score of zero means ‘‘no meaning preserved’’ and score of one hundred means ‘‘perfect meaning and grammar’’.
You can use the preceding context to evaluate the translation of the source. {src_lang} source: ‘‘{context}’’ {src_lang} preceding context: ‘‘{source}’’ {tgt_lang} translation: ‘‘{hyphothesis}’’ Score: |
A.2 Statistical Significance
We confirm the validity of our results using statistical significance tests. Particularly for testing the statistical significance of the error rate ratio between the gender groups (), we use paired bootstrap resampling (). Similarly for the statistical significance of QE ratios, we use the one-sample t-test ().
A.3 Filtering Details for Machine Translation Quality Experiments
As outlined in 5.2, to explore whether QE metrics exhibit gender bias when assessing machine-generated translations, we translate the source-feminine () and source-masculine () texts from the counterfactual subset of MTGenEval with Google Translate (GT). Then the counterfactual pairs, () and () are filtered through the following two-step process:
1.
Target gender inflection: We exclude counterfactual pairs where either of the translations do not exhibit the correct target gender inflection, as categorized by Currey et al. (2022).
2.
Translation quality consistency: We remove pairs with translations of different quality, as measured by BLEU Papineni et al. (2002). Specifically, we split translations into 5 quality categories (“Poor”,“Fair”,“Good”,“Very Good” and “Excellent”), following Lavie (2011). We retain only the pairs of instances for which both translations belong to the same quality category. Additionally, within each quality category, we confirm that BLEU does not significantly differ between gender groups using the Wilcoxon test ().
This filtering process ensures that all retained counterfactual pairs are correctly gender-inflected and of equivalent translation quality. Importantly, we retain around half of the initial instances as indicated in Table 15, most of which are of high-quality according to BLEU.
A.4 Hardware specifications
All our experiments were conducted using 2 NVIDIA RTX A6000 GPUs.
A.5 Discussion on artifacts
Appendix B Additional Results
This section reports additional figures and tables to supplement the findings in the main body.
B.1 Results on Gender Ambiguous Instances
In this part, we provide a comprehensive view of results for the ambiguous case. Specifically, Figures 5 and 6 report the raw QE scores when assessing masculine and feminine references, on MT-GenEval and GATE respectively. In addition, Figures 7 and 8 report the ratios of the QE scores on GATE and mGeNTE datasets respectively.
B.2 Results on Gender Unambiguous Instances
Additional results for Intra-sentential Case.
In this setup, gender ambiguity is resolved through contextual cues within the sentence itself. To provide a more comprehensive analysis, we include additional results for this case. Table 7 presents the complete results, including the total error rate , error rate ratio between gender groups , and QE score ratios , averaged across languages. Additionally, Tables 8 and 9 provide the per-language error rate ratios between gender groups and per-language QE score ratios , along with the corresponding results of statistical significance tests.
Additional results for Extra-sentential Case.
In this setup, gender ambiguity is resolved by contextual cues from preceding sentences. As explained in §3.1, we experiment with two variants for neural QE metrics: 1) prepending the context of the source unmodified to the candidate hypothesis, and 2) prepending the context translated to the candidate hypothesis. For GEMBA metrics we experiment with two versions for including context(see Table 6).
To provide a comprehensive view, we include the complete results. Table 10 presents the complete results, including the total error rate , the error rates for sources with masculine referents and feminine referents , and the error rate ratio between gender groups , averaged across languages. Table 11 provides the error rate ratios between gender groups , along with the corresponding results of statistical significance tests for individual languages. In addition, Table 14 summarizes the aggregated tie rates across all languages.
According to our findings, the majority of QE metrics is biased. Additionally, translating the context versus prepending it unmodified to the hypothesis, might improve accuracy but in the trade of bias.
Lastly, as indicated in Figure 9, the translation system being used to translate the context—Google Translate (◼) or NLLB (⚫)—does not affect any trend majorly. Metrics might show marginal improvements in accuracy or gender equality, however the majority of them is biased.
B.3 Results when assessing QE metrics on Automatic Translations
In § 5.2 we explore whether QE metrics demonstrate any systematic biases when applied to automatic translations. As detailed in § 5.2 we focus on the intra-sentential case, where the gender identity of the referent is resolved by the source sentence itself. We conduct our analysis on counterfactual pairs () and (), by combining the counterfactual subset of MTGenEval and outputs obtained from Google Translate (see Appendix A.3).
To provide a comprehensive view, we include the complete results on individual language pairs. Specifically, in Table 12, we report the QE score ratios, along with the corresponding statistical tests, for individual languages. In addition, Table 13 presents the per-language error rate ratios between gender groups. As shown, when QE metrics are challenged to select the correct gender inflection for an automatic translation, they demonstrate significantly higher error rates on sources with feminine referents than on their masculine counterparts. A finding inline with our observations when assessing human-written translations.
Appendix C AI Assistants
We have used Github Copilot111111https://github.com/features/copilot and Claude 3.5 Sonnet during development of our research work.
| Metric | |||
| CometKiwi 22 | 0.11 | 1.70 | 0.997 |
| CometKiwi 23 XL | 0.09 | 1.18 | 1.001 |
| CometKiwi 23 XXL | 0.07 | 0.87 | 1.003 |
| xCOMET XL | 0.10 | 1.81 | 0.988 |
| xCOMET XXL | 0.08 | 1.32 | 0.995 |
| MetricX 23 L | 0.31 | 1.25 | 0.998 |
| MetricX 23 XL | 0.12 | 1.19 | 1.000 |
| Llama 3.1 70B | 0.31 | 1.16 | 0.995 |
| Gemma 2 9B | 0.28 | 1.36 | 1.000 |
| Mistral 7B | 0.74 | 1.13 | 1.000 |
| GPT 4o | 0.16 | 1.15 | 1.001 |
| Metrics | ||||||||
| de | es | fr | it | pt | ru | hi | ar | |
| CometKiwi 22 | 2.500 | 1.909 | 2.333 | 1.773 | 0.938 | 2.500 | 0.683 | 1.000 |
| CometKiwi 23 XL | 1.143 | 1.136 | 0.500 | 1.000 | 1.333 | 2.333 | 1.297 | 0.706 |
| CometKiwi 23 XXL | 0.286 | 0.923 | 0.750 | 1.036 | 0.692 | 2.000 | 1.000 | 0.273 |
| xCOMET XL | 3.000 | 1.154 | 2.200 | 1.346 | 2.429 | 2.200 | 1.181 | 1.000 |
| xCOMET XXL | 0.857 | 1.333 | 1.143 | 1.321 | 1.111 | 2.500 | 1.426 | 0.833 |
| MetricX 23 L | 1.037 | 1.582 | 1.459 | 1.186 | 0.839 | 1.302 | 0.937 | 1.660 |
| MetricX 23 XL | 1.600 | 1.083 | 0.857 | 0.959 | 0.550 | 2.182 | 1.182 | 1.130 |
| Mistral 7B | 1.190 | 1.125 | 1.184 | 1.194 | 1.098 | 1.191 | 1.045 | 0.990 |
| Gemma 2 9B | 0.775 | 1.317 | 1.767 | 1.443 | 1.459 | 1.421 | 1.248 | 1.426 |
| Llama 3.1 70B | 0.902 | 1.265 | 1.113 | 1.161 | 1.140 | 1.185 | 1.304 | 1.210 |
| GPT 4o | 0.857 | 1.200 | 0.939 | 1.167 | 1.037 | 1.368 | 1.214 | 1.400 |
| Metrics | ||||||||
| de | es | fr | it | pt | ru | hi | ar | |
| CometKiwi 22 | 0.995 | 0.995 | 0.995 | 0.996 | 1.000 | 0.997 | 0.998 | 1.000 |
| CometKiwi 23 XL | 0.997 | 1.000 | 0.998 | 0.996 | 0.998 | 1.000 | 1.006 | 1.015 |
| CometKiwi 23 XXL | 1.005 | 0.995 | 1.006 | 0.991 | 0.999 | 1.001 | 1.006 | 1.025 |
| xCOMET XL | 0.988 | 0.975 | 0.997 | 0.979 | 0.982 | 0.991 | 0.996 | 0.995 |
| xCOMET XXL | 0.997 | 0.993 | 0.997 | 0.989 | 1.001 | 0.998 | 0.989 | 0.995 |
| MetricX 23 L | 0.999 | 0.998 | 0.998 | 0.998 | 1.000 | 0.999 | 1.000 | 0.992 |
| MetricX 23 XL | 0.998 | 0.998 | 0.998 | 1.000 | 1.001 | 1.000 | 1.002 | 1.000 |
| Llama 3.1 70B | 1.002 | 0.993 | 0.989 | 0.999 | 0.993 | 0.991 | 0.995 | 1.001 |
| Gemma 2 9B | 0.997 | 1.001 | 1.002 | 1.004 | 0.999 | 0.998 | 1.003 | 0.999 |
| Mistral 7B | 1.000 | 1.000 | 0.999 | 1.000 | 0.998 | 1.000 | 1.002 | 1.002 |
| GPT 4o | 0.999 | 1.003 | 0.998 | 1.003 | 1.000 | 0.993 | 1.000 | 1.007 |
| Metric |
CometKiwi 22 |
CometKiwi 22 |
CometKiwi 23 XL |
CometKiwi 23 XL |
CometKiwi 23 XXL |
CometKiwi 23 XXL |
xCOMET XL |
xCOMET XL |
xCOMET XXL |
xCOMET XXL |
MetricX 23 L |
MetricX 23 L |
MetricX 23 XL |
MetricX 23 XL |
Llama 3.1 70B |
Llama 3.1 70B V2 |
Gemma 2 9B |
Gemma 2 9B V2 |
Mistral 7B |
Mistral 7B V2 |
GPT- 4o |
[…]
| 0.32 | 0.10 | 0.30 | 0.09 | 0.09 | 0.08 | 0.11 | 0.09 | 0.07 | 0.09 | 0.46 | 0.28 | 0.17 | 0.24 | 0.38 | 0.38 | 0.31 | 0.34 | 0.79 | 0.75 | 0.24 | |
| 0.25 | 0.28 | 0.26 | 0.24 | 0.19 | 0.19 | 0.45 | 0.37 | 0.32 | 0.30 | 0.33 | 0.39 | 0.26 | 0.32 | 0.81 | 0.81 | 0.81 | 0.82 | 0.96 | 0.95 | 0.61 | |
| 0.84 | 6.68 | 0.92 | 5.09 | 2.20 | 4.85 | 4.53 | 7.03 | 4.97 | 5.47 | 0.73 | 1.53 | 1.54 | 1.60 | 2.20 | 2.22 | 2.71 | 2.54 | 1.22 | 1.29 | 2.64 |
| Metrics | |||||||||
| de | es | fr | it | pt | ru | hi | ar | ||
| CometKiwi 22 | 0.691 | 0.853 | 0.895 | 1.117 | 1.068 | 0.724 | 0.638 | 0.704 | |
| CometKiwi 22 | ✓ | 6.636 | 9.068 | 8.990 | 8.278 | 7.364 | 10.820 | 1.506 | 0.807 |
| CometKiwi 23 XL | 1.066 | 1.012 | 1.133 | 1.049 | 1.248 | 0.744 | 0.490 | 0.595 | |
| CometKiwi 23 XL | ✓ | 4.771 | 4.610 | 4.362 | 5.371 | 6.102 | 10.396 | 4.371 | 0.736 |
| CometKiwi 23 XXL | 2.675 | 2.597 | 2.221 | 2.161 | 2.627 | 2.027 | 2.092 | 1.162 | |
| CometKiwi 23 XXL | ✓ | 3.487 | 3.656 | 3.508 | 6.821 | 6.496 | 9.941 | 4.168 | 0.746 |
| xCOMET XL | 5.887 | 4.686 | 4.161 | 4.859 | 4.308 | 6.104 | 4.611 | 1.622 | |
| xCOMET XL | ✓ | 10.291 | 4.668 | 10.043 | 7.486 | 8.606 | 7.320 | 6.547 | 1.300 |
| xCOMET XXL | 4.461 | 6.011 | 5.759 | 6.337 | 4.848 | 3.921 | 5.675 | 2.760 | |
| xCOMET XXL | ✓ | 3.849 | 5.810 | 8.804 | 6.020 | 5.934 | 6.153 | 6.265 | 0.945 |
| MetricX 23 L | 0.559 | 0.681 | 0.650 | 0.821 | 0.837 | 0.864 | 0.823 | 0.607 | |
| MetricX 23 L | ✓ | 1.118 | 1.476 | 1.649 | 1.826 | 1.050 | 2.776 | 1.277 | 1.038 |
| MetricX 23 XL | 1.423 | 1.423 | 1.705 | 1.781 | 1.051 | 1.916 | 1.214 | 1.802 | |
| MetricX 23 XL | ✓ | 1.395 | 1.353 | 2.166 | 2.059 | 1.300 | 1.881 | 1.726 | 0.955 |
| Mistral 7B | 1.494 | 1.220 | 1.296 | 1.285 | 1.222 | 1.208 | 1.018 | 1.030 | |
| Mistral 7B V2 | 1.665 | 1.317 | 1.357 | 1.324 | 1.311 | 1.235 | 1.035 | 1.094 | |
| Gemma 2 9B | 3.129 | 3.034 | 2.869 | 2.654 | 3.458 | 2.271 | 2.127 | 2.148 | |
| Gemma 2 9B V2 | 2.972 | 2.935 | 3.029 | 2.508 | 2.861 | 2.286 | 1.959 | 1.806 | |
| Llama 3.1 70B | 2.126 | 2.597 | 2.559 | 2.397 | 2.387 | 1.844 | 2.238 | 1.447 | |
| Llama 3.1 70B V2 | 2.195 | 2.791 | 2.520 | 2.279 | 2.386 | 1.855 | 2.253 | 1.507 | |
| GPT 4o | 2.705 | 3.016 | 2.444 | 2.811 | 2.784 | 2.556 | 2.202 | 2.580 | |
| Metrics | ||||||||
| de | es | fr | it | pt | ru | hi | ar | |
| CometKiwi 22 | 0.998 | 0.997 | 0.996 | 0.996 | 0.998 | 0.996 | 0.997 | 0.997 |
| CometKiwi 23 XL | 0.999 | 0.998 | 0.996 | 1.011 | 1.004 | 0.990 | 1.006 | 0.993 |
| CometKiwi 23 XXL | 1.003 | 0.996 | 1.024 | 1.000 | 1.013 | 0.994 | 0.996 | 0.996 |
| xCOMET XL | 0.993 | 0.985 | 0.985 | 0.994 | 0.993 | 0.983 | 0.981 | 1.003 |
| xCOMET XXL | 0.999 | 0.995 | 0.994 | 0.999 | 0.999 | 0.991 | 1.005 | 0.996 |
| MetricX 23 L | 1.000 | 1.000 | 0.998 | 0.998 | 0.999 | 0.999 | 1.000 | 0.992 |
| MetricX 23 XL | 1.002 | 1.001 | 1.000 | 0.999 | 0.997 | 1.004 | 1.003 | 1.001 |
| Metrics | ||||||||
| de | es | fr | it | pt | ru | hi | ar | |
| CometKiwi 22 | 0.571 | 1.269 | 2.000 | 1.292 | 1.100 | 2.333 | 0.526 | 1.667 |
| CometKiwi 23 XL | 1.600 | 1.143 | 0.833 | 1.160 | 1.000 | 3.667 | 1.053 | 2.333 |
| CometKiwi 23 XXL | 0.600 | 1.125 | 4.000 | 1.000 | 0.667 | 1.500 | 1.030 | 3.000 |
| xCOMET XL | 2.000 | 1.238 | 4.000 | 1.304 | 2.571 | 1.800 | 0.958 | 1.778 |
| xCOMET XXL | 0.667 | 1.286 | 1.750 | 1.038 | 1.429 | 1.000 | 1.167 | 1.500 |
| MetricX 23 L | 1.086 | 1.216 | 1.115 | 1.208 | 0.973 | 1.290 | 0.966 | 1.474 |
| MetricX 23 XL | 1.429 | 1.207 | 1.286 | 1.125 | 1.400 | 2.143 | 1.071 | 1.000 |
| Metric | tr(c) | Tie rate (%) |
| MetricX 23 L | ✓ | 0.04 |
| MetricX 23 XL | 0.19 | |
| MetricX 23 XL | ✓ | 11.30 |
| xCOMET XL | 0.25 | |
| xCOMET XL | ✓ | 0.30 |
| xCOMET XXL | 1.44 | |
| xCOMET XXL | ✓ | 1.05 |
| Metric | Tie rate (%) |
| Llama 3.1 70B V1 | 55.45 |
| Llama 3.1 70B V2 | 55.01 |
| Gemma 2 9B V1 | 49.65 |
| Gemma 2 9B V2 | 51.67 |
| Mistral 7B V1 | 85.18 |
| Mistral 7B V2 | 82.48 |
| GPT 4o | 38.34 |
| Quality | # Stage 1 | # Stage 2 | |
| F | M | ||
| Excellent (50+) | 662 | 654 | 590 |
| Very Good (40-50) | 282 | 271 | 191 |
| Good (30-40) | 218 | 209 | 140 |
| Fair (20-30) | 194 | 189 | 143 |
| Poor (<20) | 149 | 163 | 114 |
| Total | 1505 | 1486 | 1178 |